# Quantum adiabatic optimization with Rydberg arrays: localization phenomena and encoding strategies

### v1

@@ V Observation of localization and optimization performance on neutral atom hardware
To efficiently probe these phenomena on the hardware, we include the tail interaction [i.e., $U_{v,u}$ in Eq. (2) beyond the blockade radius] in modeling the system, and elaborate appropriate strategies to mitigate its effect on the quantum dynamics. In particular, the tail interaction leads to site-dependent diagonal terms $\epsilon(\bm{j})$ in the two-domain-wall effective Hamiltonian in Eq. (5). In Appendix F, we discuss how we can minimize the variance of these terms (and thus recover the blockade-limit physics) by adjusting either the position of the atoms along the vertex-wires or the local detuning on the atoms. In this section, we employ the position-based mitigation strategy to counteract the tail interaction. Therefore, on the physical hardware, we implement the four configurations of the CWE setup with optimized positions shown in the left column of Fig. 9.
@@ V.1 Localized ground states
In the central column of Fig. 9, we show the expectation value of $C(\bm{j})$, evaluated for the ground state of the system computed via exact diagonalization for $\Delta=3\Omega$. Due to the contribution of states with more than two domain walls, $\langle C(\bm{j})\rangle$ slightly deviates from the perturbative ground states of Fig. 7, while preserving their fundamental localization features (in agreement with the discussion at the end of Sec. IVB). In particular, the non-zero value of $\langle C(\bm{j})\rangle$ outside of the $L$-shaped domain is due to states with more than two domain walls, and it tends to zero as $\Omega/\Delta$ decreases.
@@ V.2 Success probability
Here, we directly observe how the geometry of the CWE setup (and the associated localization physics) determines the success probability on the neutral atom hardware. We define the success probability of adiabatic optimization as the probability of measuring the solution to the encoded MWIS problem at the end of the protocol. In particular, we consider the four configurations of the CWE setup in the left column of Fig. 9 and execute the standard protocol in the bottom-left of Fig. 10. This setup encodes the simple graph of two vertices connected by an edge, and the solution to the MWIS problem is known based on the weight of the vertices. We can target this solution by properly removing one atom from the system, without resorting to the local detuning on the boundary atoms. Specifically, when $\delta_{x}>\delta_{y}$ ($\delta_{x}<\delta_{y}$), we target the $|10\rangle_{l}$ ($|01\rangle_{l}$) state by removing the last atom from the $x$ ($y$) vertex-wire.
In the top-left of Fig. 10, we show the probability $P$ of detecting the strings encoding the solution to the MWIS problem, after proper post-processing of the data (see Appendix G for details). For the maximum available time of $T=4$ µs, $P$ varies between $0.4$ and $0.6$, depending on configuration and target state. Importantly, the lowest values are observed in the cases where an unfavorable localization of the ground-state wavefunction with respect to the target solution is expected to occur. In particular, in the (b) configuration of Fig. 9, the success probability of finding the correct solution to the MWIS problem is equal to $0.59$ for the “easy” target state (i.e., $|01\rangle_{l}$) and to $0.42$ for the “hard” one (i.e., $|10\rangle_{l}$).
In Fig. 10(right), we deepen the analysis of this configuration by varying the algorithm runtime $T$ and measuring the corresponding probability $P$. As expected, at any $T$, we observe a higher success probability for the easy case than for the hard one. Notably, this difference can also be directly seen in the raw data (i.e., the data before post-processing), which are shown in the inset of the figure.
Figure 10: Observation of the success probability on the QuEra Aquila neutral atom hardware. (left) Success probability $P$ of the four configurations in the left column of Fig. 9 when executing the standard protocol (shown in the bottom) for $T=4$ µs. We target the final state by removing the final atom of the vertex-wire in the $|1\rangle_{l}$ state. (right) $P$ as a function of the algorithm runtime $T$ for the (b) configuration of Fig. 9. For each configuration and target state, we perform 1500 measurements of the final state. See Appendix G for details on the data analysis.
@@ VI Exponential improvement in the adiabatic performance
In this section, we present two different strategies to address the exponential closing of the minimum gap observed in the analysis of the CWE setup, aiming for an exponential improvement in the adiabatic performance.
The first approach leverages the effect of the geometry of the two-domain-wall Hilbert space on the minimum gap. In particular, we show that polynomial scaling along the logical protocol can be guaranteed through an unbiased extension of the first legs of the vertex-wires. The second approach exploits the domain-wall boundary condition to delocalize the ground-state wavefunction along the logical protocol. Specifically, we show that polynomial scaling can be guaranteed by modifying the structure of the CWE gadget to achieve such delocalization. In Appendix I, we present a third approach to achieve such a polynomial scaling by exploiting the control of the Rabi frequency on the atoms at the extreme of the vertex-wires.
VI.1 Modification of the domain-wall space geometry via vertex-wire extension
The first strategy relies on the dependence of the minimum gap scaling on the geometry of the two-domain-wall Hilbert space (see Fig. 7). Specifically, we notice that the gap closes exponentially with $L=N/2+1$ when the width of one (both) lead(s) is kept constant during the scaling (i.e., $\zeta_{x}$ and/or $\zeta_{y}$). In the physical system, this corresponds to keeping the first leg(s) of the two vertex-wires fixed when increasing $L$. This observation suggests a modification of the setup, where the first legs of both vertex-wires are simultaneously and unbiasedly extended by adding $2d$ atoms, chosen to be proportional to the original number of atoms per vertex-wire $N$. This increases the total number of domain-wall states per vertex-wire to $L+d$. This strategy could be straightforwardly generalized to a network by extending the initial leg of every vertex-wire.
@@ VI.1 Modification of the domain-wall space geometry via vertex-wire extension
Figure 11: Effectiveness of the vertex-wire-extension strategy. This strategy is applied (a,b) to the configuration of Fig. 4(c) and (c,d) to the configuration of Fig. 4(b). The vertex-wires (with $N$ atoms) are extended by adding $2d$ atoms to the legs before the gadget so that $L=N/2+1$ is extended to $L+d$ [see bottom-left insets in (b,d)]. (a,c) Minimum gap $\delta E_{m}$, normalized to $t_{0}=\Omega_{0}^{2}/\Delta_{0}$ (where $\Omega_{0}$ and $\Delta_{0}$ are the maximum Rabi frequency and detuning), along the logical protocol as a function of $d/L$ for a fixed $L$. (b,d) Scaling of the minimum gap for different extension ratios $d/L$ as a function of $L$.
@@ VI.2 Domain-wall delocalization via ancillary atoms
The relationship between boundary conditions and ground-state delocalization can be understood in the continuum limit. In particular, when homogeneous Neumann boundary conditions (NBC) are assigned on the whole boundary of the $L$-shaped domain, the ground state of the Helmholtz equation in Eq. (6) is the constant function, independently of the configuration.
@@
To satisfy this condition, the diagonal terms $\epsilon(\bm{j})$ need to compensate for the lack of links on the boundary sites [where #$\#\text{links}(\bm{j})<4$]. With the current structure of the gadget, this can be achieved on the logical component of the boundary by controlling the local detuning on the atoms at the extreme of the vertex-wires (setting ${\epsilon_{i,w}=\epsilon_{f,w}=t}$, with $w=x,y$). However, it cannot be achieved on the non-logical component, where $\epsilon(\bm{j})$ are forced to be identically zero.
To enforce the hopping-detuning balance condition on the non-logical component of the boundary, we propose to modify the structure of the CWE gadget by introducing properly designed ancillary atoms, as depicted in Fig. 12(a) (see details in Appendix H). In this way, we extend the boundary of the two-domain-wall Hilbert space by creating additional domain-wall states, whose diagonal terms can be controlled via local detuning on the ancillae. The effectiveness of this strategy is demonstrated in Fig. 12(b), where we employ this modification of the gadget to impose Eq. (9) at half-protocol. We observe a polynomial scaling of the minimum gap for all configurations and target states.
Figure 12: Modification of the crossing-with-edge gadget via the introduction of ancillary atoms. (a) Six ancillae are added to the gadget structure, with detuning $\Delta+\epsilon_{A}$ on $A_{w}$, $\Delta$ on $A_{Nw}$, and $\epsilon_{B}$ on $B_{w}$, where $w=x,y$. This graph can be drawn explicitly as a unit-disk graph. This modification effectively extends the non-logical component of the two-domain-wall Hilbert space boundary by introducing additional virtual sites depicted in orange and blue. Here, the diagonal terms are given by the detuning on the ancillae: $\epsilon_{A}$ on the orange sites and $\epsilon_{A}+\epsilon_{B}$ on the blue sites. (b) Minimum gap scaling as a function of $L=N/2+1$ (with $N$ being the number of atoms per vertex-wire) for the four configurations of Fig. 7 when modifying the gadget structure as in (a) and imposing Eq. (9) at $\tau=T/2$ in the logical protocol. For the configuration of Fig. 7(b), the data corresponding to the two final ground states $|01\rangle_{l}$ and $|10\rangle_{l}$ are shown in orange and blue, respectively.
@@ VII Conclusion
In this work, we analyzed the quantum properties of the Rydberg encoding scheme for maximum weighted independent set problems on arbitrarily connected graphs introduced in Ref. [19]. In particular, we focused the analysis on its fundamental building blocks — two vertex-wires intersecting via a gadget — for various geometrical configurations.
We quantified the adiabatic performance of these systems by looking at the minimum gap scaling with system size along two representative adiabatic protocols. Our analysis shows that, for some geometrical configurations, a quantum mechanical effect leads to an unfavorable localization of the ground-state wavefunction with respect to the solution to the optimization problem, potentially resulting in an exponential closing of the minimum gap. While we discussed this phenomenon in the simplest instances of the encoding scheme in Ref. [19], we expect this effect to also arise in more complex instances and potentially in other encoding schemes.
Our work shows, through a concrete example, that a careful design of encoding strategies for quantum adiabatic optimization requires going beyond a mere classical analysis of the target Hamiltonian. It also necessitates a detailed consideration of the interplay between the encoding strategy and the quantum mechanical effects that occur during the adiabatic algorithm. In the last part of this work, we gave an example of this approach, showing how the encoding strategy of Ref. [19] can be adapted to avoid quantum bottlenecks.
@@ Acknowledgment
We thank Giuliano Giudici for valuable discussions on developing the intersecting tensor network method. We also thank Sheng-Tao Wang, Milan Kornjaca, and John Long for their helpful discussions. We acknowledge Amazon Web Services for providing credits to access the Aquila neutral atom machine. This work is supported by the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101113690 (PASQuanS2.1), the ERC Starting grant QARA (Grant No. 101041435), the EU-QUANTERA project TNiSQ (N-6001), the Austrian Science Fund (FWF) (Grant No. DOI 10.55776/COE1), the DARPA ONISQ program (Grant No. W911NF2010021), the US Department of Energy (DOE Quantum Systems Accelerator Center, Grant No. DE-AC02-05CH11231), the Center for Ultracold Atoms (an NSF Physics Frontiers Center), and the National Science Foundation. SN acknowledges funding from the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie Grant No. 101059826 (ETNA4Ryd) and support from the Quantum Computing and Simulation Center (QCSC) of Padova University.
@@ Appendix A Intersecting tensor network
Here, we illustrate the intersecting tensor network (ITN) method designed for simulating two vertex-wires intersecting via a gadget [see Fig. 1(c,d)].
Let us start by considering the case of a single vertex-wire, where it is possible to apply the MPS-DMRG method. This method consists of approximating the state of the system with a matrix product state (MPS) ansatz and variationally searching for the ground state via the DMRG algorithm [84].
To construct the matrix product operator (MPO) for the Rydberg Hamiltonian within the blockade-constraint subspace, we employ an approach that can be easily extended from the vertex-wire to the gadget setups. Let us start from the Rydberg Hamiltonian in the generic form of Eq. (2). In the blockade limit, it can be reformulated into the following “PXP” form: $H_{\rm Rydberg}=\sum_{v}\mathcal{P}h_{v}\mathcal{P}$, where $h_{v}=~{}\Omega\sigma^{x}_{v}-\Delta_{v}n_{v}$, and $\mathcal{P}$ is the projector into the blockade-constraint subspace, which comprises of all states with no neighboring atoms in $|1\rangle$. Practically, within the sum, we can substitute the global projector $\mathcal{P}$ with the local ones $\mathcal{P}_{v}\equiv\prod_{u\in\mathcal{B}(v)}p_{u}$, where $p_{u}=|0\rangle_{u}\langle 0|$ and $\mathcal{B}(v)$ denotes the set of atoms blockaded with the atom $v$. For a single vertex-wire, we can express the Rydberg Hamiltonian exactly in the form of an MPO, where we associate to each atom $v$ the following tensor:
@@
We now consider the case of two intersecting vertex-wires. Here, the state of the system can be expressed as a generalization of the MPS for a vertex-wire, as shown in Fig. 14(a). Here, the atoms on the vertex-wires are represented by tensors with two virtual indices and one physical index. The atoms at the intersection of the two vertex-wires — the gadget atoms for the crossing setup and the gadget atoms plus two other atoms for the CWE setup — are grouped into a gadget tensor with four virtual indices and one physical index. In representing the state, we truncate the dimension of the virtual indices to a maximum bond dimension of $D$. The dimension of the physical indices $d$ is equal to $2$ for the vertex-wire, $5$ for the crossing gadget tensor, and $8$ for the CWE gadget tensor. Importantly, for this structure, it is feasible to employ the single-site update DMRG algorithm to variationally search for the ground states and low-excited states [64, 65, 66, 85].
@@
From this representation of the state, we can then construct the corresponding MPO form of the Rydberg Hamiltonian for the gadget setups. As for the single vertex-wire, the tensor corresponding to a vertex-wire atom $v$ is given by $W^{[v]}$ in Eq. (10), while a larger tensor $W^{[\mathrm{gad}]}$ corresponds to the gadget atoms. This tensor has four virtual indices of dimension $4$ and two physical indices of dimension $d=5$ for the crossing setup and $d=8$ for the CWE setup. We graphically represent its non-zero entries in Fig. 13.
@@
Figure 13: Graphical representation of MPOs for the vertex-wire, the crossing, and the crossing-with-edge (CWE) gadget in Fig. 1. We follow the notation of Ref. [86], where the numbers indicate the position of the operators within the element. For instance, $2\rightarrow\boxed{h}\rightarrow 3$ denotes that operator $h$ is located at the $2$nd row and $3$rd column of $W$ in Eq. (10). For the gadgets, the placing of the operators inside the boxes corresponds to the position of the atoms in the physical structure. The physical indices are omitted for simplicity.
@@
Figure 14: Intersecting-tensor-network method for the gadget setups. (a) Graphical representation of the state as a product of tensors. Tensors of two virtual indices (in black) and one physical index (in gray) correspond to the vertex-wire atoms. The tensor depicted in orange of four virtual indices (in black) and one physical index (in gray) corresponds to the atoms at the intersection of the vertex-wires, shown in the inset. (b,c) Analysis of the convergence of the DMRG method for the crossing setup with vertex-wires (of $N$ atoms) intersecting at the middle ($m_{x}=m_{y}=N/2$), $\Delta=1.3\Omega$, and $\Delta_{g}=3\Delta$. (b) Difference between the ground-state energy obtained via exact diagonalization ($E_{\rm exact}$) and DMRG ($E_{D}$) as a function of the maximum bond dimension $D$, for $N=10$. (c) Difference between the ground-state energy obtained using DMRG with bound dimension $D<32$ and $D=32$, for $N=102$.
@@
In Fig. 14, we test the convergence of the algorithm for increasing $D$ in the case of the crossing setup. In (b), we show the difference between the ground-state energy computed using DMRG (as $D$ varies) and that computed via exact diagonalization, for a small system of $N=10$ atoms per vertex-wire. In (c), we compare it with the ground-state energy computed using DMRG with the largest accessible bound dimension $D=32$ for a large system of $N=102$ atoms per vertex-wire. The largest accessible bound dimension is mainly limited by the dimension of the gadget tensor ($dD^{4}$). From this analysis, we see that a bond dimension of $D=16$ allows us to achieve sufficiently small errors. In this work, we set $D=16$. Similar considerations hold for the CWE setup.
Appendix B Effective Hamiltonian for vertex-wires

### v2

@@ V Observation of localization and optimization performance on neutral atom hardware
To efficiently probe these phenomena on the hardware, we include the tail interaction [i.e., $U_{v,u}$ in Eq. (2) beyond the blockade radius] in modeling the system, and elaborate appropriate strategies to mitigate its effect on the quantum dynamics. In particular, the tail interaction leads to site-dependent diagonal terms $\epsilon(\bm{j})$ in the two-domain-wall effective Hamiltonian in Eq. (5). In Appendix F, we discuss how we can minimize the variance of these terms (and thus recover the blockade-limit physics) by adjusting either the position of the atoms along the vertex wires or the local detuning on the atoms. In this section, we employ the position-based mitigation strategy to counteract the tail interaction. Therefore, on the physical hardware, we implement the four configurations of the CWE setup with optimized positions shown in the left column of Fig. 9.
@@ V.1 Localized ground states
In the central column of Fig. 9, we show the expectation value of $C(\bm{j})$, evaluated for the ground state of the system computed via exact diagonalization for $\Delta=3\Omega$. Due to the contribution of states with more than two domain walls, $\langle C(\bm{j})\rangle$ slightly deviates from the perturbative ground states of Fig. 7, while preserving their fundamental localization features (in agreement with the discussion at the end of Sec. IVB). In particular, the nonzero value of $\langle C(\bm{j})\rangle$ outside of the $L$-shaped domain is due to states with more than two domain walls, and it tends to zero as $\Omega/\Delta$ decreases.
@@ V.2 Success probability
Here, we directly observe how the geometry of the CWE setup (and the associated localization physics) determines the success probability on the neutral atom hardware. We define the success probability of adiabatic optimization as the probability of measuring the solution to the encoded MWIS problem at the end of the protocol. In particular, we consider the four configurations of the CWE setup in the left column of Fig. 9 and execute the standard protocol in the bottom left of Fig. 10. This setup encodes the simple graph of two vertices connected by an edge, and the solution to the MWIS problem is known based on the weight of the vertices. We can target this solution by properly removing one atom from the system, without resorting to the local detuning on the boundary atoms. Specifically, when $\delta_{x}>\delta_{y}$ ($\delta_{x}<\delta_{y}$), we target the $|10\rangle_{l}$ ($|01\rangle_{l}$) state by removing the last atom from the $x$ ($y$) vertex wire.
In the top-left of Fig. 10, we show the probability $P$ of detecting the strings encoding the solution to the MWIS problem, after proper postprocessing of the data (see Appendix G for details). For the maximum available time of $T=4$ µs, $P$ varies between $0.4$ and $0.6$, depending on configuration and target state. Importantly, the lowest values are observed in the cases where an unfavorable localization of the ground-state wavefunction with respect to the target solution is expected to occur. In particular, in the (b) configuration of Fig. 9, the success probability of finding the correct solution to the MWIS problem is equal to $0.59$ for the “easy” target state (i.e., $|01\rangle_{l}$) and to $0.42$ for the “hard” one (i.e., $|10\rangle_{l}$).
In Fig. 10(right), we deepen the analysis of this configuration by varying the algorithm runtime $T$ and measuring the corresponding probability $P$. As expected, at any $T$, we observe a higher success probability for the easy case than for the hard one. Notably, this difference can also be directly seen in the raw data (i.e., the data before postprocessing), which are shown in the inset of the figure.
Figure 10: Observation of the success probability on the QuEra Aquila neutral atom hardware. (Left) Success probability $P$ of the four configurations in the left column of Fig. 9 when executing the standard protocol (shown in the bottom) for $T=4$ µs. We target the final state by removing the final atom of the vertex wire in the $|1\rangle_{l}$ state. (Right) $P$ as a function of the algorithm runtime $T$ for the (b) configuration of Fig. 9. For each configuration and target state, we perform 1500 measurements of the final state. See Appendix G for details on the data analysis.
@@ VI Exponential improvement in the adiabatic performance
In this section, we present two different strategies to avoid the exponential closing of the minimum gap observed in the analysis of the CWE setup. This results in an exponential improvement in the adiabatic performance and suggests a ‘quantum-aware’ modification of the encoding strategy of Ref. [21].
The first approach leverages the effect of the geometry of the two-domain-wall Hilbert space on the minimum gap. In particular, we show that polynomial scaling along the logical protocol can be guaranteed through an unbiased extension of the first legs of the vertex wires. The second approach exploits the domain-wall boundary condition to delocalize the ground-state wavefunction along the logical protocol. Specifically, we show that polynomial scaling can be guaranteed by modifying the structure of the CWE gadget to achieve such delocalization. In Appendix I, we present a third approach to achieve such a polynomial scaling by exploiting the control of the Rabi frequency on the atoms at the extreme of the vertex wires.
VI.1 Modification of the domain-wall space geometry via vertex wire extension
The first strategy relies on the dependence of the minimum gap scaling on the geometry of the two-domain-wall Hilbert space (see Fig. 7). Specifically, we notice that the gap closes exponentially with $L=N/2+1$ when the width of one (both) lead(s) is kept constant during the scaling (i.e., $\zeta_{x}$ and/or $\zeta_{y}$). In the physical system, this corresponds to keeping the first leg(s) of the two vertex wires fixed when increasing $L$. This observation suggests a modification of the setup, where the first legs of both vertex wires are simultaneously and unbiasedly extended by adding $2d$ atoms, chosen to be proportional to the original number of atoms per vertex wire $N$. This increases the total number of domain-wall states per vertex wire to $L+d$.
@@ VI.1 Modification of the domain-wall space geometry via vertex wire extension
This strategy could be generalized to a network of gadgets encoding an arbitrarily connected graph of $|V|$ vertices [see Fig. 1(a) and Ref. [21]] by adding $2d$ atoms at the beginning of each wire. Assuming that a similar analysis holds in this general case, the optimal choice of $d/L\simeq 0.5$, suggested by Fig. 11(c), would increase the atom-number overhead of the encoding scheme from the original scaling of $4|V|^{2}$ to $4|V|^{2}+2d|V|\simeq 6|V|^{2}$.
Figure 11: Effectiveness of the vertex wire-extension strategy. This strategy is applied (a,b) to the configuration of Fig. 4(c) and (c,d) to the configuration of Fig. 4(b). The vertex wires (with $N$ atoms) are extended by adding $2d$ atoms to the legs before the gadget so that $L=N/2+1$ is extended to $L+d$ [see bottom-left insets in (b,d)]. (a,c) Minimum gap $\delta E_{m}$, normalized to $t_{0}=\Omega_{0}^{2}/\Delta_{0}$ (where $\Omega_{0}$ and $\Delta_{0}$ are the maximum Rabi frequency and detuning), along the logical protocol as a function of $d/L$ for a fixed $L$. (b,d) Scaling of the minimum gap for different extension ratios $d/L$ as a function of $L$.
@@ VI.2 Domain-wall delocalization via ancillary atoms
The relationship between boundary conditions and ground-state delocalization can be understood in the continuum limit. In particular, when homogeneous Neumann boundary conditions (NBCs) are assigned on the whole boundary of the $L$-shaped domain, the ground state of the Helmholtz equation in Eq. (6) is the constant function, independently of the configuration.
@@
To satisfy this condition, the diagonal terms $\epsilon(\bm{j})$ need to compensate for the lack of links on the boundary sites [where #$\#\text{links}(\bm{j})<4$]. With the current structure of the gadget, this can be achieved on the logical component of the boundary by controlling the local detuning on the atoms at the extreme of the vertex wires (setting ${\epsilon_{i,w}=\epsilon_{f,w}=t}$, with $w=x,y$). However, it cannot be achieved on the nonlogical component, where $\epsilon(\bm{j})$ are forced to be identically zero.
To enforce the hopping-detuning balance condition on the nonlogical component of the boundary, we propose to modify the structure of the CWE gadget by introducing properly designed ancillary atoms, as depicted in Fig. 12(a) (see details in Appendix H). In this way, we extend the boundary of the two-domain-wall Hilbert space by creating additional domain-wall states, whose diagonal terms can be controlled via local detuning on the ancillae. The effectiveness of this strategy is demonstrated in Fig. 12(b), where we employ this modification of the gadget to impose Eq. (9) at half-protocol. We observe a polynomial scaling of the minimum gap for all configurations and target states.
Figure 12: Modification of the crossing-with-edge gadget via the introduction of ancillary atoms. (a) Six ancillae are added to the gadget structure, with detuning $\Delta+\epsilon_{A}$ on $A_{w}$, $\Delta$ on $A_{Nw}$, and $\epsilon_{B}$ on $B_{w}$, where $w=x,y$. This graph can be drawn explicitly as a unit-disk graph. This modification effectively extends the nonlogical component of the two-domain-wall Hilbert space boundary by introducing additional virtual sites depicted in orange and blue. Here, the diagonal terms are given by the detuning on the ancillae: $\epsilon_{A}$ on the orange sites and $\epsilon_{A}+\epsilon_{B}$ on the blue sites. (b) Minimum gap scaling as a function of $L=N/2+1$ (with $N$ being the number of atoms per vertex wire) for the four configurations of Fig. 7 when modifying the gadget structure as in (a) and imposing Eq. (9) at $\tau=T/2$ in the logical protocol. For the configuration of Fig. 7(b), the data corresponding to the two final ground states $|01\rangle_{l}$ and $|10\rangle_{l}$ are shown in orange and blue, respectively.
This strategy could also be applied to a network of gadgets. In this case, since the minimal unit-disk realization of the modified CWE gadget in Fig. 12(a) requires 17 atoms, the atom-number overhead of the encoding scheme would increase from the original scaling of $4|V|^{2}$ to $8.5|V|^{2}$.
@@ VII Conclusion
In this work, we analyzed the quantum properties of the Rydberg encoding scheme for maximum weighted independent set problems on arbitrarily connected graphs introduced in Ref. [21]. In particular, we focused the analysis on its fundamental building blocks — two vertex wires intersecting via a gadget — for various geometrical configurations.
We quantified the adiabatic performance of these systems by looking at the minimum gap scaling with system size along two representative adiabatic protocols. Our analysis shows that the interplay between geometry and domain-wall kinetic energy can lead to an unfavorable localization of the ground-state wavefunction with respect to the solution to the optimization problem, and to an exponential closing of the minimum gap. Finally, we proposed two strategies to address this localization and provide exponential improvement in the adiabatic performance. Our findings are summarized in Table. 1.
Setup Standard Logical Logical/improv.
Crossing poly poly poly
CWE—(a) exp poly poly
CWE—(b) poly/exp poly/exp poly
CWE—(c) unclear exp poly
CWE—(d) poly poly poly
Table 1: Summary of the minimum gap scaling for the two setups along the standard protocol, the logical protocol, and the logical protocol with improvement strategies. For the CWE setup, we refer to the geometrical configurations of Figs. 4 and 7. The asterisk denotes that the scaling depends on the target state.
While we discussed the localization phenomenon in the simplest instances of the encoding scheme in Ref. [21], we expect this effect to also arise in more complex instances and in other encoding schemes. For example, in the case of a network of $|V|$ intersecting vertex wires, the $L$-shaped domain-wall space is generalized to a truncated $|V|$-dimensional hypercube. In such a domain, we anticipate the presence of analogous localization phenomena. While it is possible to extend our improvement strategies to this general case, quantifying their effectiveness for large networks remains an open problem. In addition, it would be interesting to consider alternative strategies designed to prevent quantum optimization algorithms from being trapped in suboptimal solutions, such as diabatic sweeps or counteradiabatic driving [86, 87, 88, 74, 89].
In summary, our work shows, through a concrete example, that a careful design of encoding strategies for quantum adiabatic optimization requires going beyond a mere classical analysis of the target Hamiltonian. It also necessitates a detailed consideration of the interplay between the encoding strategy and the quantum mechanical effects that occur during the adiabatic algorithm. Our work provides an example of such a ‘quantum-aware’ approach, showing how the encoding strategy of Ref. [21] can be adapted to avoid quantum bottlenecks.
@@ Acknowledgment
We thank G. Giudici for valuable discussions on developing the intersecting tensor network method. We also thank S.-T. Wang, M. Kornjaca, and J. Long for their helpful discussions. We acknowledge Amazon Web Services for providing credits to access the Aquila neutral atom machine. This work is supported by the European Union’s Horizon Europe research and innovation program under Grant Agreement No. 101113690 (PASQuanS2.1), the ERC Starting grant QARA (Grant No. 101041435), the EU-QUANTERA project TNiSQ (N-6001), the Austrian Science Fund (FWF) (Grant No. DOI 10.55776/COE1), the DARPA ONISQ program (Grant No. W911NF2010021), the US Department of Energy (DOE Quantum Systems Accelerator Center, Grant No. DE-AC02-05CH11231), the Center for Ultracold Atoms (an NSF Physics Frontiers Center), and the National Science Foundation. SN acknowledges funding from the European Union’s Horizon Europe research and innovation program under the Marie Skłodowska-Curie Grant No. 101059826 (ETNA4Ryd) and support from the Quantum Computing and Simulation Center (QCSC) of Padova University.
@@ Appendix A Intersecting tensor network
Here, we illustrate the intersecting tensor network (ITN) method designed for simulating two vertex wires intersecting via a gadget [see Fig. 1(c,d)].
Let us start by considering the case of a single vertex wire, where it is possible to apply the MPS-DMRG method. This method consists of approximating the state of the system with a matrix product state (MPS) ansatz and variationally searching for the ground state via the DMRG algorithm [90].
To construct the matrix product operator (MPO) for the Rydberg Hamiltonian within the blockade-constraint subspace, we employ an approach that can be easily extended from the vertex wire to the gadget setups. Let us start from the Rydberg Hamiltonian in the generic form of Eq. (2). In the blockade limit, it can be reformulated into the following “PXP” form: $H_{\rm Rydberg}=\sum_{v}\mathcal{P}h_{v}\mathcal{P}$, where $h_{v}=~{}\Omega\sigma^{x}_{v}-\Delta_{v}n_{v}$, and $\mathcal{P}$ is the projector into the blockade-constraint subspace, which comprises of all states with no neighboring atoms in $|1\rangle$. Practically, within the sum, we can substitute the global projector $\mathcal{P}$ with the local ones $\mathcal{P}_{v}\equiv\prod_{u\in\mathcal{B}(v)}p_{u}$, where $p_{u}=|0\rangle_{u}\langle 0|$ and $\mathcal{B}(v)$ denotes the set of atoms blockaded with the atom $v$. For a single vertex wire, we can express the Rydberg Hamiltonian exactly in the form of an MPO, where we associate to each atom $v$ the following tensor:
@@
We now consider the case of two intersecting vertex wires. Here, the state of the system can be expressed as a generalization of the MPS for a vertex wire, as shown in Fig. 14(a). Here, the atoms on the vertex wires are represented by tensors with two virtual indices and one physical index. The atoms at the intersection of the two vertex wires — the gadget atoms for the crossing setup and the gadget atoms plus two other atoms for the CWE setup — are grouped into a gadget tensor with four virtual indices and one physical index. In representing the state, we truncate the dimension of the virtual indices to a maximum bond dimension of $D$. The dimension of the physical indices $d$ is equal to $2$ for the vertex wire, $5$ for the crossing gadget tensor, and $8$ for the CWE gadget tensor. Importantly, for this structure, it is feasible to employ the single-site update DMRG algorithm to variationally search for the ground states and low-excited states [66, 67, 68, 91].
@@
From this representation of the state, we can then construct the corresponding MPO form of the Rydberg Hamiltonian for the gadget setups. As for the single vertex wire, the tensor corresponding to a vertex wire atom $v$ is given by $W^{[v]}$ in Eq. (10), while a larger tensor $W^{[\mathrm{gad}]}$ corresponds to the gadget atoms. This tensor has four virtual indices of dimension $4$ and two physical indices of dimension $d=5$ for the crossing setup and $d=8$ for the CWE setup. We graphically represent its nonzero entries in Fig. 13.
@@
Figure 13: Graphical representation of MPOs for the vertex wire, the crossing, and the CWE gadget in Fig. 1. We follow the notation of Ref. [92], where the numbers indicate the position of the operators within the element. For instance, $2\rightarrow\boxed{h}\rightarrow 3$ denotes that operator $h$ is located at the second row and third column of $W$ in Eq. (10). For the gadgets, the placing of the operators inside the boxes corresponds to the position of the atoms in the physical structure. The physical indices are omitted for simplicity.
@@
Figure 14: Intersecting-tensor-network method for the gadget setups. (a) Graphical representation of the state as a product of tensors. Tensors of two virtual indices (in black) and one physical index (in gray) correspond to the vertex-wire atoms. The tensor depicted in orange of four virtual indices (in black) and one physical index (in gray) corresponds to the atoms at the intersection of the vertex wires, shown in the inset. (b,c) Analysis of the convergence of the DMRG method for the crossing setup with vertex wires (of $N$ atoms) intersecting at the middle ($m_{x}=m_{y}=N/2$), $\Delta=1.3\Omega$, and $\Delta_{g}=3\Delta$. (b) Difference between the ground-state energy obtained via exact diagonalization ($E_{\rm exact}$) and DMRG ($E_{D}$) as a function of the maximum bond dimension $D$, for $N=10$. (c) Difference between the ground-state energy obtained using DMRG with bound dimension $D<32$ and $D=32$, for $N=102$.
@@
In Fig. 14, we test the convergence of the algorithm for increasing $D$ in the case of the crossing setup. In (b), we show the difference between the ground-state energy computed using DMRG (as $D$ varies) and that computed via exact diagonalization, for a small system of $N=10$ atoms per vertex wire. In (c), we compare it with the ground-state energy computed using DMRG with the largest accessible bound dimension $D=32$ for a large system of $N=102$ atoms per vertex wire. The largest accessible bound dimension is mainly limited by the dimension of the gadget tensor ($dD^{4}$). From this analysis, we see that a bond dimension of $D=16$ allows us to achieve sufficiently small errors. In this work, we set $D=16$. Similar considerations hold for the CWE setup.
Appendix B Effective Hamiltonian for vertex wires